Given a population of 5000 score with the mean and . How many scores are:



1
Expert's answer
2021-05-07T13:08:52-0400

Total problem:

given the population of 5000 scores with mean = 86 and standard deviation = 10.

How many score are:

a. between 86 and 96?

B. between 96 and 106?

c. Below 56%

a.

"P(86<X<96) = P(X<96) -P(X<86) \\

= P(Z< \frac{96-86}{10}) -P(Z< \frac{86-86}{10}) \\

= P(Z<1) -P(Z<0) \\

= 0.8413 - 0.5 \\

= 0.3413 \\

N = 0.3413 \times 5000 = 1706.5"

b.

"P(96<X<106) = P(X<106) -P(X<96) \\

= P(Z< \frac{106-86}{10}) -P(Z< \frac{96-86}{10}) \\

= P(Z<2) -P(Z<1) \\

= 0.9772-0.8413 \\

= 0.1359 \\

N = 0.1359 \times 5000 = 679.5"

c. "N = 0.56 \times 5000 = 2800"